6 oz of copper which costs $4/oz were combined with 2 oz of tin which costs 8/oz. Find the cost per oz of the mixture. please put steps on how to do it.
Answer by stanbon(75887)

6 oz of copper which costs $4/oz were combined with 2 oz of tin which costs 8/oz. Find the cost per oz of the mixture. please put steps on how to do it.
-------
Total value = 6*4+2*8 = 24+16 = $40.00
-----
Value per ounce = 40/(6+2) = $5.00
======================
Cheers,
Stan H.
==========